# More with Less: An Empirical Study of Turn-Control Strategies for Efficient Coding Agents

### v1

@@ 6.1. Optimizing the Prompt
The most direct approach to reducing cost is optimizing the prompt. Prompt compression methods shorten the input while preserving key semantic content (Hu et al., 2025; Jiang et al., 2023). These generally fall into two categories: extractive techniques that prune redundant tokens based on information-theoretic metrics, and abstractive techniques that use a smaller model to generate a summary (Li et al., 2024). More advanced methods leverage the model’s internal states (Fei et al., 2025) or reinforcement learning (Hu et al., 2025; Jiang et al., 2023) to select the most salient information.
A more holistic approach is Retrieval-Augmented Generation (RAG), a form of context engineering where an agent retrieves relevant information from an external knowledge base (e.g., documentation or code) to augment its prompt before generation (Shinn et al., 2023; Lewis et al., 2020; GitHub, 2025; Su et al., 2024). By improving the prompt’s signal-to-noise ratio, RAG directly addresses the "lost in the middle" problem, where models struggle to utilize information buried in long contexts (Liu et al., 2023).
@@ 6.2. Re-architecting the Agent
A second category of optimizations moves beyond the prompt to re-design the agent’s fundamental architecture. Agentic Plan Caching stores and reuses structured plan templates for semantically similar tasks. This allows the system to bypass the expensive, multi-turn planning process for known problems, invoking a full planner LLM only for novel tasks. This has been shown to reduce serving costs by 46.62% while maintaining 96.67% of optimal performance (Zhang et al., 2025).
Codified Prompting, as seen in the CodeAgents framework (Yang et al., 2025), replaces verbose natural language communication between agent components with a structured, typed pseudocode. This makes the agent’s reasoning more precise and token-efficient, reducing input tokens by 55%–87% and output tokens by 41%–70%.
Hierarchical and Multi-Agent Systems decompose complex tasks, delegating sub-tasks to multiple, specialized agents (Wu et al., 2024; CrewAI, 2024; LangChain, 2024). This enables strategies like cost-effective cascades, where cheaper models handle most work and expensive models are reserved for critical steps, achieving cost reductions up to 94.2% (Gandhi et al., 2024).
@@ 6.3. Constraining the Reasoning Process
The research most adjacent to our work involves direct constraints on the agent’s reasoning process. Token Budgeting explicitly instructs the LLM to limit the length of its generated response in a given turn. The TALE framework, for example, dynamically estimates a "reasonable" token budget and injects it into the prompt, reducing token costs by an average of 67% (Han et al., 2024). Other work has explored fine-tuning models to self-estimate budgets (Li et al., 2025). A key finding from this research is the "Token Elasticity" phenomenon: an overly aggressive budget can backfire, causing the model to generate an even longer response (Han et al., 2024).
In practice, many systems also use a Static Step/Turn Limit (e.g., max_turns) as a brute-force safeguard against infinite loops, as seen in frameworks like SWE-bench (Jimenez et al., 2023; Cognition AI, 2024). However, this is not an intelligent strategy, as a fixed limit lacks the adaptability for tasks of varying complexity.
@@ 7. Conclusion
The code and related experimental data in this paper are accessibleon this page: https://anonymous.4open.science/r/icse26_turn_control-778D/README.md

### v2

@@ 4.2. Case Study 2
4.3. Stochasticity of Case Study
We conduct supplementary repeated experiments (10 runs for each case study) to verify the stochasticity and robustness of them, as illustrated in Table 6
Table 6. Repeated Experiments for Each Case Study
Case Study Configuration Result in Paper Repeated Runs Success Rate 95% Clopper-Pearson
Case 1 75th-percentile Solved 8 Solved 80% [0.44, 0.97]
Unlimited-turn Unsolved 4 Solved 40% [0.12, 0.74]
Case 2 Dynamic 50->75 Solved 4 Solved 40% [0.12, 0.74]
Fixed 75 Unsolved 2 Solved 20% [0.03, 0.56]
In two completely independent cases, our proposed strategies demonstrated a strong effect of doubling the success rate (from 40% to 80% in Case 1, and from 20% to 40% in Case 2). Due to the small sample size (n=10), a result of the high cost of these experiments, these differences do not reach statistical significance at the $p<0.05$ level. However, both cases together reveal a clear and robust positive trend. For example, in Case 1, the lower bound of the constrained strategy’s confidence interval (0.44) is already higher than the point estimate of the unlimited-turn strategy (0.40), which clearly indicates a strong trend of superior performance.
These repeated experiments greatly enhance the credibility of our case studies and the robustness of our method. They indicate that the performance improvements we observed are not accidental, but rather a systematic and repeatable positive effect of our proposed turn-control strategies.
@@ 5.1. Stochasticity of LLMs
5.2. LLM Selection
Another threat to validity lies in the selection of proprietary LLMs. Our model selection focused on state-of-the-art models from the top commercial providers, namely OpenAI, Google, and Anthropic, to ensure our findings are relevant to the most widely used, production-grade agentic systems. At the time of our experiments, on OpenRouter, the popular model service provider, nine of the top ten most-used models are closed-source models while the only open-source model is free to use, indicating that for paid workloads where cost-efficiency is critical, the market overwhelmingly relies on closed-source solutions.
The decision for Sonnet over Opus was data-driven. While Opus is a top performer, Sonnet is far more widely deployed. For instance, recent OpenRouter data shows Sonnet’s minimum daily usage (33.4 billion tokens) is over five times higher than Opus’s peak (6.35 billion). For a study on practical cost-efficiency, selecting the more heavily utilized model ensures our findings have a significantly greater impact on real-world agentic workloads.
@@ 6.1. Optimizing the Prompt
The most direct approach to reducing cost is optimizing the prompt. Prompt compression methods shorten the input while preserving key semantic content (hu2025dynamic; jiang2023llmlingua). These generally fall into two categories: extractive techniques that prune redundant tokens based on information-theoretic metrics, and abstractive techniques that use a smaller model to generate a summary (li2024prompt). More advanced methods leverage the model’s internal states (fei2025efficient) or reinforcement learning (hu2025dynamic; jiang2023llmlingua) to select the most salient information.
A more holistic approach is Retrieval-Augmented Generation (RAG), a form of context engineering where an agent retrieves relevant information from an external knowledge base (e.g., documentation or code) to augment its prompt before generation (shinn2023reflexion; lewis2020rag; github2024rag; su2024evor). By improving the prompt’s signal-to-noise ratio, RAG directly addresses the "lost in the middle" problem, where models struggle to utilize information buried in long contexts (liu2023lost).
@@ 6.2. Re-architecting the Agent
A second category of optimizations moves beyond the prompt to re-design the agent’s fundamental architecture. Agentic Plan Caching stores and reuses structured plan templates for semantically similar tasks. This allows the system to bypass the expensive, multi-turn planning process for known problems, invoking a full planner LLM only for novel tasks. This has been shown to reduce serving costs by 46.62% while maintaining 96.67% of optimal performance (zhang2025cost).
Codified Prompting, as seen in the CodeAgents framework (yang2025codeagents), replaces verbose natural language communication between agent components with a structured, typed pseudocode. This makes the agent’s reasoning more precise and token-efficient, reducing input tokens by 55%–87% and output tokens by 41%–70%.
Hierarchical and Multi-Agent Systems decompose complex tasks, delegating sub-tasks to multiple, specialized agents (wu2024autogen; crewai2024; langchain2024memory). This enables strategies like cost-effective cascades, where cheaper models handle most work and expensive models are reserved for critical steps, achieving cost reductions up to 94.2% (gandhi2024budgetmlagent).
@@ 6.3. Constraining the Reasoning Process
The research most adjacent to our work involves direct constraints on the agent’s reasoning process. Token Budgeting explicitly instructs the LLM to limit the length of its generated response in a given turn. The TALE framework, for example, dynamically estimates a "reasonable" token budget and injects it into the prompt, reducing token costs by an average of 67% (han2024token). Other work has explored fine-tuning models to self-estimate budgets (li2025selfbudgeter). A key finding from this research is the "Token Elasticity" phenomenon: an overly aggressive budget can backfire, causing the model to generate an even longer response (han2024token).
In practice, many systems also use a Static Step/Turn Limit (e.g., max_turns) as a brute-force safeguard against unlimited loops, as seen in frameworks like SWE-bench (jimenez2023swebench; cognition2024swebench). However, this is not an intelligent strategy, as a fixed limit lacks the adaptability for tasks of varying complexity.
@@ 7. Conclusion
The code and related experimental data in this paper are accessible on this page: https://doi.org/10.6084/m9.figshare.30705524